OpenStudy (anonymous):

How do I solve for x? log3(x+9)+log3(x-9)=1

OpenStudy (mathteacher1729):

The left side of the equation can be re-written as \(\log_3(x^2-81)\). Now solve \(\log_3(x^2-81) = 1\) using the fact that \(\log_3(3) = 1\)

OpenStudy (anonymous):

write the sum of the logs as the log of the product\[\log_{3}[(x+9)(x-9)]=1 \]then use the def of log to write\[3^1=(x+9)(x-9)\]now solve the polynomial\[x^2-81=3\]\[x=\pm \sqrt{84}\]=pm 2 sqrt{21}

OpenStudy (anonymous):

posted prematurely\[= \pm 2 \sqrt{21}\]both of these have to be checked for domain\[\sqrt{84} \approx 9.16\]so the negatove one must be discarded; the solution set is \[\left\{ 2 \sqrt{21} \right\}\]

OpenStudy (anonymous):

ok...thanks for clearing that part up

OpenStudy (anonymous):

ur welcome; I accidentally posted and would have got to it all at once; i seem to be all thumbs tonight :{

OpenStudy (anonymous):

been there....thanks again :D

OpenStudy (anonymous):

:})